A stack of playing cards contains 4 jacks, 5 queens, 3 kings, and 3 aces. Two cards will be randomly selected from the stack. What is the probability that a queen is chosen and replaced, and then a queen is chosen again?
Answer by greenestamps(13200)



There are 5 queens out of a total of 15 cards.

Since the first queen is replaced before selecting the second card, the probability of getting a queen on each selection is 5/15 = 1/3.

The probability of getting a queen both times is then (1/3)(1/3) = 1/9.

ANSWER: 1/9